OpenStudy (anonymous):

A firecracker shoots up from a hill 165 feet high with an initial speed of 110 feet per second. Using the formula H(t) = -16t2 + vt + s, determine how long it will take the firecracker to hit the ground

OpenStudy (binary3i):

put H(t)=0 and s=165 and solve the quadratic equation.

OpenStudy (jit4won):

Total distance covered in vertical direction = - 165 ft Initial velocity in vertical direction = 0 acceleration = -32ft/s^2 using S = ut + 1/2 at^2 -165 = 1/2 x -32 x t^2 t^2 = 165/16 you can calculate time now